# Hybrid-z: Enhancing Kilo-Degree Survey bright galaxy sample photometric redshifts with deep learning

### v1

@@ 5.2 Validation of Hybrid-z across external spectroscopic samples
Table 2: Statistics of photometric redshift performance for KiDS-Bright cross-matched with various spectroscopic samples.
@@
    Number of galaxies in the cross-matched sample.
@@
In addition to the standard evaluation of Hybrid-z performance from Sec. 5.1 done on test samples statistically consistent with the training data (as both were randomly sampled from input GAMA equatorial catalogs), we have done a further validation of our photo-$z$s. Still using GAMA-Equatorial for training, we checked Hybrid-z predictions for a number of spectroscopic samples overlapping with the KiDS DR4 area and covering appropriate magnitude ranges and sky areas to provide sufficient statistics. These include GAMA Equatorial and G23 fields (Driver et al., 2022), 2dFGRS (Colless et al., 2001), SDSS DR16 (Ahumada et al., 2020), and 2dFLenS (Blake et al., 2016). Among these, GAMA G23 and 2dFLenS are disjoint with our GAMA-Equatorial training set, as they cover only the southern patch of KiDS (at $\delta<-25^{\circ}$) [^6] A very small fraction ($\sim 1\%$) of 2dFLenS is located in KiDS-N, but the overlap with GAMA-Equatorial fields is negligible ($\sim 200$ sources).. SDSS and 2dFGRS do have common objects with GAMA in its equatorial patches, they however extend to the much wider KiDS area, so their overlap with GAMA-eq is also small.
@@
In Table 2 we present the statistics derived from cross-matching the above-mentioned datasets with the full KiDS-Bright ‘clean’ sample. For each spectroscopic dataset, we compare the performance of Hybrid-z with previous ANNz2 results from B21. For each of these test sets, Hybrid-z performs considerably better in terms of scatter than ANNz2, while generally the former displays larger bias than the latter. What is worth noticing is the very small SMAD of Hybrid-z ($\sim 0.012$) for the SDSS DR16 cross-match. This dataset, due to the $r<20$ flux limit of KiDS-Bright, includes galaxies from the SDSS main sample (flux-limited to $r<17.77$), as well as a subset of BOSS LOWZ and CMASS, dominated by luminous red galaxies. This combination leads to the photo-z statistics for such a SDSS $\times$ KiDS-Bright cross-match being overall better than for the general red galaxies specified in Table 1. However, the same is not the case for ANNz2, which performs only marginally better for such a mixture of SDSS galaxies than it did for the general red galaxy selection.
@@
Among the datasets included in Table 2, GAMA G23 and 2dFLenS are genuinely ‘blind’ test samples, as they are separated on-sky by many degrees from the area where the GAMA-Equatorial training sets are. Photo-$z$ statistics derived from these catalogs should be therefore robust against possible overfitting thanks to their independence from the training data. While one should remember that neither G23 nor 2dFLenS are as complete and flux-limited samples as KiDS-Bright and GAMA-Equatorial, it is reassuring to find that for both of the former Hybrid-z gives considerable improvement over ANNz2, namely reduction in scatter by respectively 17% and 20%.
@@ 5.3 Application to the KiDS-DR4 Bright sample
The Hybrid-z model trained on the full GAMA equatorial data, as discussed above, provides satisfactory results when tested on various overlapping spectroscopic datasets. However, applying it directly to the KiDS-Bright sample introduces artifacts in the final photo-$z$ distribution, which are related to the specific properties of the GAMA training set. Namely, GAMA sky coverage of 180 deg2 is small enough that the survey is affected by cosmic variance, manifesting itself by enhanced effects of the cosmic web, such as voids and filaments (e.g. Eardley et al., 2015). This results in significant ‘peaks’ and ‘dips’ in the GAMA redshift distribution, which would average out for larger sky area. This then imprints onto our photo-$z$ predictions. Namely, training the Hybrid-z model directly on the full GAMA dataset leads to $dN/dz_{\mathrm{phot}}$ of KiDS-Bright which mimics some of the LSS-related properties of GAMA $dN/dz_{\mathrm{spec}}$. Furthermore, GAMA is complete to a brighter magnitude than the KiDS-Bright selection (Jalan et al., 2024), which will lead to some extrapolation of model predictions at the faint end.
@@
As discussed in previous papers (e.g. Bilicki et al., 2016, B21), the redshift distribution of GAMA has strong features, and in particular, there is an under-abundance of galaxies at $z_{\mathrm{spec}}\sim 0.25$. This redshift is close to the median of the sample, which is where ML models typically work optimally. If we then train on such a specific distribution, our model becomes biased towards this input $dN/dz$, which results in a ‘dip’ at $z_{\mathrm{phot}}\sim 0.25$ and two peaks below and above this value
(see panel (a) of Fig. 7). As photo-zs dilute the structures in the radial direction, for a sample covering $\sim 1000$ deg2 such as ours, this kind of strong features in its $dN/dz_{\mathrm{phot}}$ are not physical but rather originate from ‘redshift focusing’ of the model. Indeed, this behavior was not observed in B21, where photo-$z$s were of worse precision than here. Our interpretation is that the improved sensitivity of our new model reveals these artifacts, indicating that training data specifics now significantly influence model behavior.
@@
The second effect we observe is related to the mismatch between the training and target photometric datasets at the faint end. As quantified in the recent paper by Jalan et al. (2024), the GAMA-equatorial sample becomes considerably incomplete with respect to KiDS-Bright at $r\gtrsim 19.5$ mag. For a supervised ML model like ours, this leads to extrapolation resulting from the so-called covariate shift. ML models tend to perform well when the test and training data share a well-matched feature space and distribution of the target quantity. When this distribution shifts or a covariate shift occurs, model performance can be adversely affected (Y et al., 2019). In our case, this affects the predictions at the faintest magnitudes of the sample. In particular, as was already the case for ANNz2, many galaxies with $r>19.5$ are assigned $z_{\mathrm{phot}}>0.35$ where training data is sparser. For Hybrid-z, training directly on the full GAMA dataset introduces additionally a new peak at $z_{\mathrm{phot}}\sim 0.38$, revealing a redshift focusing effect in our model’s predictions as evident from panel (a) of Fig. 7).
@@
In order to solve the above-discussed issues, for the final training of the full-sample photo-zs, we decided to subsample the spectroscopic redshifts from the GAMA equatorial dataset in such a way as to smooth out the peaks and dips originally present. The subsampling was done in such a way to not affect the color-redshift relation but instead provide a ‘smoother’ (more regular) input redshift distribution in GAMA without the strong features discussed above. We create this smoothed subsample of training data by iteratively adjusting the distribution of redshift values to achieve a more uniform histogram. Starting with an initial histogram of true $z$ values, bins with counts that significantly exceed their neighboring bins are identified as spikes, based on a decreasing threshold. In each iteration, if a bin exceeds the count of its neighbors by this threshold, its count is reduced to either the average of its neighbors or its original count, whichever is lower. Randomly selected data points are then drawn from each bin according to the adjusted counts, preserving the dataset’s structure but smoothing out extreme values. This iterative process results in a more evenly distributed subsample, which is useful for downstream tasks, by reducing overrepresented regions. Finally, the resulting ‘KiDS ID’ values and their corresponding $z$ values for this subsample are output, with a calculation of the subsample size as a percentage of the original dataset. This gave us an output of 118k galaxies from the GAMA training set (about $66\%$of the original one) and their redshift distribution is shown as red bars in panel (b) in Fig.7. Using these smoothened data we retrained the Hybrid-z model and then applied it to the full $r<20$ mag dataset of $\sim 1.2$ million KiDS DR4 galaxies. Among these, about 996k have the flag ‘masked=0’ indicating their usefulness for science (see B21 for details).
@@
To summarize, the Hybrid-z model, trained on the full GAMA equatorial dataset, performs well on overlapping spectroscopic surveys, but introduces artifacts when applied directly to the KiDS-Bright sample, due to the specific properties of the GAMA training set. To mitigate these problems, a smoothed subsample of the training data was created to achieve a more uniform redshift distribution, which was then used to retrain the model. As a result, we obtain photo-z predictions for the full KiDS DR4 Bright galaxy sample, which displays improved performance over previous derivations and also gives a generally artifact-free redshift distribution.
@@
Figure 7: Photo-$z$ distribution of the KiDS-DR4 bright sample ($r<20$ mag) as predicted by ANNz2 (blue, B21), compared with photo-$z$ estimates from Hybrid-z (this work) trained on spec-$z$ distribution of (a) KiDS x GAMA bright sources (original training sample); (b) smoothened KiDS x GAMA bright sources.
@@ 6 Conclusions and future prospects
This work presents the first deep-learning photometric redshift (photo-$z$) derivations for the flux-limited KiDS-Bright galaxy sample, with the selection threshold $r<20$ mag (Bilicki et al., 2021). Previously, photo-$z$s for this catalog were estimated using ‘shallow’ learning methods, specifically the ANNz2 neural network package (Sadeh et al., 2016). Our new model, called Hybrid-z, is built upon recent studies, including Li et al. (2022), where deep learning was applied for photo-$z$s in a deeper KiDS galaxy sample, and Treyer et al. (2024), who used a similar training set as us, to obtain CNN-based redshifts for an SDSS-selected catalog with the same flux limit.
@@
We have built and tested a DL model for photo-$z$ derivation, called Hybrid-z, that uses 4-band KiDS images ($ugri$) processed by CNNs, which is combined with 9-band magnitudes from KiDS+VIKING, processed by an ONNs. Rather than simply averaging the two networks’ outputs, we concatenate their outputs before the final three dense layers. Such an approach is inspired by the previous works of e.g. Li et al. (2022) or Henghes et al. (2022) and it yields significantly improved photo-$z$ performance compared to the previous ANNz2 derivations of B21 where 9-band KiDS+VIKING magnitudes were used. The Hybrid-z model reduces the scatter (SMAD) of $\Delta z$ by 20% as compared to ANNz2 for the same test samples. This holds for multiple spectroscopic test datasets, of which some can be considered entirely ‘blind’ in terms of being fully disjoint with the training data. When tested on the fiducial ‘clean’ KiDS-Bright sample, Hybrid-z achieves SMAD$(\Delta z)$ of approximately $0.014(1+z)$, representing a clear improvement over the prior ANNz2 results giving $\sim 0.018(1+z)$, while maintaining the same minimal bias in $\delta z$ of at most a few times $10^{-4}$.
@@
The Hybrid-z model shows even greater improvement in photo-$z$ precision if we separate out blue galaxies from the KiDS-Bright sample. For these objects, Hybrid-z reduces the SMAD of $\Delta z$ by 22% as compared to ANNz2, and the resulting scatter is then at the same level as the latter approach was previously providing for red galaxies. This of course means that at the same time, the photo-$z$ performance for red galaxies improves less than on average (by 17% in SMAD) after adding deep learning as compared to the ordinary ‘shallow’ networks. This is consistent with expectations: CNNs can leverage the detailed, varied features in blue, typically spiral, galaxies more effectively than in the smoother, elliptical red galaxies. In a flux-limited sample at low redshift, as ours, blue galaxies are more abundant than the red ones, so such a property of the photo-$z$ model is very useful to improve the overall quality of the derivations. These advancements pave the way for more refined astrophysical and cosmological analyses using KiDS-Bright data, such as of the stellar-to-halo-mass relation (B21) or multi-probe analyses (Dvornik et al., 2023).
@@
Our new photo-$z$ model trained on the GAMA equatorial dataset performs reliably on KiDS-overlapping spectroscopic data, but introduces artifacts in the redshift distribution when directly applied to the KiDS-Bright sample. These artifacts stem from GAMA’s limited sky coverage and cosmic variance, which affect the resulting $dN/dz_{\mathrm{phot}}$ and introduce specific patterns mimicking GAMA characteristics, such as a notable dip at $z_{\mathrm{phot}}\sim 0.25$, and extrapolation issues at faint and of $r>19.5$ mag. To overcome these limitations, we created a smoothed subsample of the GAMA training data, reducing sharp peaks and dips in its redshift distribution and achieving a more uniform $dN/dz_{\mathrm{spec}}$ for final model training. This approach mitigates artifacts in the output $dN/dz_{\mathrm{phot}}$ enhancing the reliability of our photo-$z$ estimates. These improved photo-$z$s for the KiDS-DR4 Bright Sample ($\sim$1.2M galaxies) are available publicly at https://kids.strw.leidenuniv.nl/DR4/brightsample.php.
@@
In this work, our DL analysis focused on the 4-band KiDS optical images, while the full 9-band KiDS+VIKING photometry was used only in the form of magnitudes. Moving forward, we plan to build an extended DL model incorporating images from all nine bands, from KiDS $u$ to VIKING $K_{s}$, once the NIR coadds are available – currently being prepared for the 4MOST WAVES target selection (Driver et al., 2019). We anticipate that this expansion will improve CNN-based photo-$z$ precision, mirroring the gains seen from KiDS DR3 $ugri$ (Bilicki et al., 2018) to 9-band inclusion in DR4 (Bilicki et al., 2021). We plan to apply such an enhanced Hybrid-z model to the final KiDS Data Release 5 (Wright et al., 2024) for state-of-the-art bright-end photometric redshifts. Additional improvement in DR5 is expected thanks to the second $i$-band pass which deepens effective imaging and should help further refine our photo-$z$ estimates. Last but not least, our approach holds promise for even deeper imaging applications in the forthcoming Legacy Survey of Space and Time (LSST Science Collaboration et al., 2009), paving the way for robust photo-$z$s in future large-scale sky surveys.

### v2

@@ 5.2 Validation of Hybrid-z across external spectroscopic samples
Table 2: Statistics of photometric redshift performance for KiDS-Bright crossmatched with various spectroscopic samples.
@@
    Number of galaxies in the crossmatched sample.
@@
In addition to the standard evaluation of Hybrid-z performance from Sec. 5.1 done on test samples statistically consistent with the training data (as both were randomly sampled from input GAMA equatorial catalogs), we have done a further validation of our photo-$z$s. Still using GAMA-Equatorial for training, we checked Hybrid-z predictions for a number of spectroscopic samples overlapping with the KiDS DR4 area and covering appropriate magnitude ranges and sky areas to provide sufficient statistics. These include GAMA Equatorial and G23 fields (Driver et al., 2022), 2dFGRS (Colless et al., 2001), SDSS DR16 (Ahumada et al., 2020), and 2dFLenS (Blake et al., 2016). Among these, GAMA G23 and 2dFLenS are disjoint with our GAMA-Equatorial training set, as they cover only the southern patch of KiDS (at $\delta<-25^{\circ}$) [^6] A very small fraction ($\sim 1\%$) of 2dFLenS is located in KiDS-N, but the overlap with GAMA-Equatorial fields is negligible ($\sim 200$ sources).. The surveys SDSS and 2dFGRS do have common objects with GAMA in its equatorial patches. However, they extend to the much wider KiDS area, so their overlap with GAMA-eq is also small.
In Table 2 we present the statistics derived from crossmatching the above-mentioned datasets with the full KiDS-Bright ‘clean’ sample. For each spectroscopic dataset, we compare the performance of Hybrid-z with previous ANNz2 results from B21. For each of these test sets, Hybrid-z performs considerably better in terms of scatter than ANNz2, while generally the former displays larger bias than the latter. This indicates that our new model loses some of the accuracy that ANNz2 had while gaining in precision (typical ML trade-off). We note, however, that in all the inspected cases, for Hybrid-z the mean biases in $\delta z$ or $\Delta z$ are much smaller than the scatter, meaning that the model is still highly accurate in its photo-z predictions. In the future, we plan to inspect this further with KiDS DR5 (Wright et al., 2024) 9-band imaging and extending the training with, for example, DESI DR1 (DESI Collaboration et al., 2025), to see if we can minimize both precision and accuracy at the same time.
@@
What is worth noticing is the very small SMAD of Hybrid-z ($\sim 0.012$) for the SDSS DR16 crossmatch. This dataset, due to the $r<20$ flux limit of KiDS-Bright, includes galaxies from the SDSS main sample (flux-limited to $r<17.77$), as well as a subset of BOSS LOWZ and CMASS, dominated by luminous red galaxies. This combination leads to the photo-z statistics for such a SDSS $\times$ KiDS-Bright crossmatch being overall better than for the general red galaxies specified in Table 1. However, the same is not the case for ANNz2, which performs only marginally better for such a mixture of SDSS galaxies than it did for the general red galaxy selection.
@@
Among the datasets included in Table 2, GAMA G23 and 2dFLenS are genuinely ‘blind’ test samples, as they are separated on-sky by many degrees from the area where the GAMA-Equatorial training sets are. Photo-$z$ statistics derived from these catalogs should therefore be robust against possible overfitting thanks to their independence from the training data. While one should remember that neither G23 nor 2dFLenS are as complete and flux-limited samples as KiDS-Bright and GAMA-Equatorial, it is reassuring to find that for both of the former Hybrid-z gives considerable improvement over ANNz2, namely reduction in scatter by respectively 17% and 20%.
@@ 5.3 Application to the KiDS-DR4 Bright sample
The Hybrid-z model trained on the full GAMA equatorial data, as discussed above, provides satisfactory results when tested on various overlapping spectroscopic datasets. However, applying it directly to the KiDS-Bright sample introduces artifacts in the final photo-$z$ distribution, which we believe are related to the specific properties of the GAMA training set. Namely, GAMA sky coverage of 180 deg2 is small enough that the survey is affected by cosmic variance, manifesting itself by enhanced effects of the cosmic web, such as voids and filaments (e.g., Eardley et al., 2015). This results in significant ‘peaks’ and ‘dips’ in the GAMA redshift distribution, which should average out for larger sky area. However, these kind of features are still imprinted onto our photo-$z$ predictions. Namely, training the Hybrid-z model directly on the full GAMA dataset leads to $dN/dz_{\mathrm{phot}}$ of KiDS-Bright which mimics some of the LSS-related properties of GAMA $dN/dz_{\mathrm{spec}}$.
@@
As discussed in previous papers (e.g., Bilicki et al., 2016, B21), the redshift distribution of GAMA has strong features, and in particular, there is an under-abundance of galaxies at $z_{\mathrm{spec}}\sim 0.25$. This redshift is close to the median of the sample, which is where ML models typically work optimally. If we then train on such a specific distribution, our model becomes biased toward this input $dN/dz$, which results in a ‘dip’ at $z_{\mathrm{phot}}\sim 0.25$ and two peaks below and above this value (see panel (a) of Fig. 7). As photo-zs dilute the structures in the radial direction, for a sample covering $\sim 1000$ deg2 such as ours, this kind of strong features in its $dN/dz_{\mathrm{phot}}$ are unlikely to be physical but rather originate from ‘redshift focusing’ of the model. However, this behavior was not observed in B21, where photo-$z$s were derived with ANNz2.
Our interpretation is that our new model is more sensitive to such strong features in the training data and this needs to be mitigated. One possibility would be to include training sets covering more of the sky, and hence they would be less affected by cosmic variance. This will be possible thanks to, for example, DESI. [^7] DESI Data Release 1 (DESI Collaboration et al., 2025) was released after this work had been completed; hence, we do not include it here. Another option is to train several models, with for instance different random seeds and/or architectures, and appropriately combine their outputs into the final prediction. Such an approach is implemented in ANNz2, but it uses much less computationally demanding ONNs. Employing a similar framework with DL would be beyond our scope of research. Below we propose another mitigation strategy consisting of appropriately resampling (”smoothing”) the available training set.
@@
The second effect we observe is related to the mismatch between the training and target photometric datasets at the faint end. Namely, GAMA is complete to a brighter magnitude than the KiDS-Bright selection, which will lead to some extrapolation of model predictions at the faint end. As quantified in the recent paper by Jalan et al. (2024), the GAMA-equatorial sample becomes considerably incomplete with respect to KiDS-Bright at $r\gtrsim 19.5$ mag. For a supervised ML model such as ours, this leads to extrapolation resulting from the so-called covariate shift. ML models tend to perform well when the test and training data share a well-matched feature space and distribution of the target quantity. When this distribution shifts or a covariate shift occurs, model performance can be adversely affected (Y et al., 2019). In our case, this affects the predictions at the faintest magnitudes of the sample. In particular, as was already the case for ANNz2, many galaxies with $r>19.5$ are assigned $z_{\mathrm{phot}}>0.35$ where training data is sparser. For Hybrid-z, training directly on the full GAMA dataset introduces additionally a new peak at $z_{\mathrm{phot}}\sim 0.38$, revealing a redshift focusing effect in our model’s predictions as evident from panel (a) of Fig. 7).
@@
In order to solve the above-discussed issues, for the final training of the full-sample photo-zs, we decided to subsample the spectroscopic redshifts from the GAMA equatorial dataset in such a way as to smooth out the peaks and dips originally present. The subsampling was done in such a way to not affect the color-redshift relation but instead provide a ‘smoother’ (more regular) input redshift distribution in GAMA without the strong features discussed above. We create this smoothed subsample of training data by iteratively adjusting the distribution of redshift values to achieve a more uniform histogram. Starting with an initial histogram of true $z$ values, bins with counts that significantly exceed their neighboring bins are identified as spikes, based on a decreasing threshold. In each iteration, if a bin exceeds the count of its neighbors by this threshold, its count is reduced to either the average of its neighbors or its original count, whichever is lower. Randomly selected data points are then drawn from each bin according to the adjusted counts, preserving the dataset’s structure but smoothing out extreme values. This iterative process results in a more evenly distributed subsample, which is useful for downstream tasks, by reducing overrepresented regions. Finally, the resulting ‘KiDS ID’ values and their corresponding $z$ values for this subsample are output, with a calculation of the subsample size as a percentage of the original dataset. This gave us an output of 118k galaxies from the GAMA training set (about $66\%$of the original one) and their redshift distribution is shown as red bars in panel (b) in Fig.7. Using these smoothed data we retrained the Hybrid-z model and then applied it to the full $r<20$ mag dataset of $\sim 1.2$ million KiDS DR4 galaxies. Among these, about 996k have the flag ‘masked=0’ indicating their usefulness for science (see B21 for details).
@@
To summarize, the Hybrid-z model, trained on the full GAMA equatorial dataset, performs well on overlapping spectroscopic surveys, but introduces artifacts when applied directly to the KiDS-Bright sample, due to the specific properties of the GAMA training set. To mitigate these problems, a smoothed subsample of the training data was created to achieve a more uniform redshift distribution, which was then used to retrain the model. As a result, we obtain photo-z predictions for the full KiDS-DR4 Bright galaxy sample, which displays improved performance over previous derivations and also gives a generally artifact-free redshift distribution.
@@
Figure 7: Photo-$z$ distribution of the KiDS-DR4 Bright sample ($r<20$ mag) as predicted by ANNz2 (blue, B21) compared with photo-$z$ estimates from Hybrid-z (this work) trained on the spec-$z$ distribution of (a) KiDS x GAMA bright sources (original training sample) (b) smoothed KiDS x GAMA bright sources.
@@ 6 Conclusions and future prospects
This work presents the first DL photometric redshift (photo-$z$) derivations for the flux-limited KiDS-Bright galaxy sample with the selection threshold $r<20$ mag (Bilicki et al., 2021). Previously, photo-$z$s for this catalog were estimated using ”shallow” learning methods, specifically the ANNz2 neural network package (Sadeh et al., 2016). Our new model, Hybrid-z, is built on recent studies, including Li et al. (2022), where DL was applied for photo-$z$s in a deeper KiDS galaxy sample, and Treyer et al. (2024), who used a similar training set as us to obtain CNN-based redshifts for an SDSS-selected catalog with the same flux limit.
@@
We built and tested a DL model for photo-$z$ derivation called Hybrid-z that uses four-band KiDS images ($ugri$) processed by CNNs, which are combined with nine-band magnitudes from KiDS+VIKING, processed by an ONN. Rather than simply averaging the outputs of the two networks, we concatenated their outputs before the final three dense layers. This approach was inspired by the previous works of, for example, Li et al. (2022) and Henghes et al. (2022), and it yielded significantly improved photo-$z$ performance compared to the previous ANNz2 derivations of B21, where nine-band KiDS+VIKING magnitudes were used. The Hybrid-z model reduces the scatter (SMAD) of $\Delta z$ by 20% as compared to ANNz2 for the same test samples. This is true for multiple spectroscopic test datasets, of which some can be considered entirely ”blind” in terms of being fully disjointed with the training data. When tested on the fiducial ’clean” KiDS-Bright sample, Hybrid-z achieves a SMAD$(\Delta z)$ of approximately $0.014(1+z)$, representing a clear improvement over the prior ANNz2 results giving $\sim 0.018(1+z)$ while maintaining the same minimal bias in $\delta z$ of at most a few times $10^{-4}$.
@@
The Hybrid-z model shows even greater improvement in photo-$z$ precision when we separate out blue galaxies from the KiDS-Bright sample. For these objects, Hybrid-z reduces the SMAD of $\Delta z$ by 22% as compared to ANNz2, resulting in a scatter comparable to that which ANNz2 attained for red galaxies. This of course, means that at the same time, the photo-$z$ performance for red galaxies improves less than on average (by 17% in SMAD) after adding DL as compared to the ordinary ”shallow” networks. This is consistent with expectations, as CNNs can leverage the detailed varied features in blue (typically spiral) galaxies more effectively than in the smoother elliptical red galaxies. In a flux-limited sample at low redshift, as ours, blue galaxies are more abundant than red galaxies, so such a property of the photo-$z$ model is very useful to improve the overall quality of the derivations. These advancements pave the way for more refined astrophysical and cosmological analyses using KiDS-Bright data, such as of the stellar-to-halo-mass relation (B21) or multi-probe analyses (Dvornik et al., 2023).
@@
Our new photo-$z$ model trained on the GAMA equatorial dataset performs reliably on KiDS overlapping spectroscopic data, but it introduces artifacts in the redshift distribution when directly applied to the KiDS-Bright sample. These artifacts stem from GAMA’s limited sky coverage and cosmic variance, which affect the resulting $dN/dz_{\mathrm{phot}}$ and introduce specific patterns that mimic GAMA characteristics, such as a notable dip at $z_{\mathrm{phot}}\sim 0.25$, and extrapolation issues at faint magnitudes and where $r>19.5$ mag. To overcome these limitations, we created a smoothed subsample of the GAMA training data by reducing sharp peaks and dips in its redshift distribution and achieving a more uniform $dN/dz_{\mathrm{spec}}$ for final model training. This approach mitigates artifacts in the output $dN/dz_{\mathrm{phot}}$, thus enhancing the reliability of our photo-$z$ estimates. These improved photo-$z$s for the KiDS-DR4 Bright sample ($\sim$1.2M galaxies) are available publicly at https://kids.strw.leidenuniv.nl/DR4/brightsample.php.
@@
In this work, our DL analysis focused on the four-band KiDS optical images, while the full nine-band KiDS+VIKING photometry was used only in the form of magnitudes. Moving forward, we plan to build an extended DL model incorporating images from all nine bands, from KiDS $u$ to VIKING $K_{s}$, once the NIR coadds are available; currently, they are being prepared for the 4MOST WAVES target selection (Driver et al., 2019). We anticipate that this expansion will improve CNN-based photo-$z$ precision and will mirror the gains seen from KiDS DR3 $ugri$ (Bilicki et al., 2018) to nine-band inclusion in DR4 (Bilicki et al., 2021). We plan to apply such an enhanced Hybrid-z model to the final KiDS Data Release 5 (Wright et al., 2024) for state-of-the-art bright-end photometric redshifts. Additional improvement in DR5 is expected thanks to the second $i$-band pass, which deepens effective imaging and should help further refine our photo-$z$ estimates. Last but not least, our approach holds promise for even deeper imaging applications in the forthcoming Legacy Survey of Space and Time (LSST Science Collaboration et al., 2009), paving the way for robust photo-$z$s in future large-scale sky surveys.
Data availability
The photometric redshift catalog based on the Hybrid-z DL model, containing redshifts for over 1.2 million galaxies in the KiDS-Bright DR4 sample, is available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr(130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/698/A276.